A ball projected from ground vertically upward is at same height at time t1 and t2 .the speed of projection of ball is.
 

Grade:11

2 Answers

Shobhit
7 years ago
Since particle is atsame height at time t1&t2, therefore,t1 + t2 = total time of flightt1 + t2 =2TT=t1+t2/2AlsoT=u/g from (v=u+at and v being 0 at maximum height and g is negative)Therefore u/g=T1+t2/2Or U=1/2g(t1+t2)
Soup tik Choudhury
5 years ago
 
Since particle is at same height at time t1&t2,
Therefore,
(t1 + t2)= total time of flight
(t1 + t2) =2T
T=(t1+t2)/2
AlsoT=u/g from (v=u+at and v being 0 at maximum height and g is negative)
Therefore,
u/g=(t1+t2)/2
u=g(t1+t2)/2